### v2

# Measuring Response Latencies Under Asymmetric Routing

Abstract.
Latency is a key indicator of Internet service performance. Continuously tracking the latency of client requests enables service operators to quickly identify bottlenecks, perform adaptive resource allocation or routing, and mitigate attacks. Passively measuring response latency at intermediate vantage points is attractive because it provides insight into the experience of real clients without requiring client instrumentation or incurring probing overhead.
We argue that existing passive measurement techniques have not caught up with recent trends in service deployments, specifically the increasing prevalence of obscured or encrypted transport headers and the use of asymmetric routing by design. Existing methods are inapplicable, inaccurate, or inefficient.
This paper presents Pirate, a passive approach for measuring response latencies when only client-to-server traffic is visible, even when transport headers are invisible. Pirate estimates the time gap between causal pairs—two requests such that the response to the first triggered the second—as a proxy for the client-side response latency. Our experiments with a realistic web application show that Pirate can accurately estimate the response latencies measured at the client application layer. A Pirate-enhanced layer-4 load balancer (with DSR) cuts tail latencies by 37%.
copyright: noneconference: ; 2026
1. Introduction
Latency is a key indicator of the performance and quality of interactive Internet services. For developers of such services, it is well known that smaller client-visible latencies drive better user engagement (114; 75; R. Kohavi, R. M. Henne, and D. Sommerfield (2007); 77; 80). Given its primacy, accurate latency measurements can feed important decisions in the design and adaptation of networked systems. For example, operators of some large content delivery services use the latency between content servers and users to determine which servers to redirect users to (Krishnan et al., 2009; Szymaniak et al., 2008). Autonomous Systems (ASes) on the wide-area Internet may use high or variable latencies experienced by transiting connections to identify pathologies such as persistent link congestion (Dhamdhere et al., 2018; Sundaresan et al., 2017; Choi et al., 2007; Sommers et al., 2007) or interdomain route hijack (107; 125; A. Arnbak and S. Goldberg (2015)), and take corrective actions to fix routing configurations or provision capacity (Lin et al., 2025). Within a data center, server latency may be used to implement performance-optimized replica selection in load balancers (Barbette et al., 2020; Gandhi and Narayana, 2025) or remote procedure call (RPC) clients (Wydrowski et al., 2024; Suresh et al., 2015).
Continuous measurement of client-visible latency is crucial, as latency may change over the lifetime of a connection (Høiland-Jørgensen et al., 2016; Aikat et al., 2003; Ding and Rabinovich, 2015; Barford and Crovella, 1999). For example, time-varying bursts and packet losses in the network (Ghabashneh et al., 2022; Canel et al., 2024), or server variability due to noisy neighbors, load, and resource scheduling (Lo et al., 2015; Gan et al., 2019; Cai et al., 2021; Barford and Crovella, 1999), can significantly change the latency perceived by the same client connection over time.
Consequently, the community has developed several techniques and systems for continuous latency measurement. Broadly, active approaches send explicit probes to observe latency, for example, by running ICMP pings between servers (Guo et al., 2015). Passive approaches observe latency at strategic locations that can be controlled (Zhu et al., 2015; Szymaniak et al., 2008; Yu et al., 2011), possibly at intermediate vantage points outside the client and the server. Passive approaches do not require instrumentation or self-reporting from clients, which may be untrustworthy or difficult to change (Szymaniak et al., 2008). If passive measurement is possible, it can observe real and representative client connections (Barford and Crovella, 1999), avoiding the typical downsides of active approaches that incur compute and network resources for probing (Choffnes and Bustamante, 2009; Kompella et al., 2009).
This Paper. We are interested in passively and continuously measuring response latencies, which we define as the time between when an application-layer request is sent and the last byte of the response is delivered to the client application, i.e., the per-object time-to-last-byte. For RPCs within data centers, response latency corresponds to the RPC completion time at the application layer (Li et al., 2014; Kalia et al., 2019). For web-based applications, the response latencies of specific objects (e.g., those first painted or largest on the user’s screen) are highly correlated with several quality-of-experience metrics (E. Bocchi, L. De Cicco, and D. Rossi (2016); 22; 66; 5). As we elaborate in §2.1, to apply to a broad range of scenarios, we seek techniques to passively measure response latency while meeting the following additional requirements:
  1. (1)
    handle missing or encrypted transport headers;
  2. (2)
    handle routing asymmetry;
  3. (3)
    generalize across transport dynamics;
  4. (4)
    support efficient deployment on software middleboxes.
Application-layer response latency is most closely related to the transport-layer round-trip time (RTT), whose passive measurement is widely studied, e.g.(Shawn Ostermann, ; Ghasemi et al., 2017; Sengupta et al., 2022; Veal et al., 2005; Jaiswal et al., 2004; Carra et al., 2010; Ding and Rabinovich, 2015; Jiang and Dovrolis, 2002). However, response latency is distinct from the transport-layer RTT, since the server may return a transport-layer acknowledgment well before the full application-layer response. As far as we are aware, no existing RTT measurement technique meets the additional requirements above, and we know of no passive techniques for directly measuring response latency (§2.2).
Our Key Ideas. This paper uses three key ideas to meet our measurement goals (§3).
Our first idea is to leverage the closed-loop nature of Internet protocols and applications. When application requests depend on the contents of prior responses—for example, a web object embeds other objects—the reception of a previous response generates the subsequent request. Flow-controlled applications (transports) mandate that new requests (packets) be transmitted only after prior responses (acknowledgments) arrive. Such closed-loop packet transmission behavior enables estimating the response latency by proxy: the vantage point can measure the time delay between a request and a subsequent request that was triggered by the reception of the response to the first request. We call the latter request a causally-triggered request, and the pair of requests a causal pair. Many latency-sensitive applications exhibit cross-request dependencies (e.g., web (Netravali et al., 2016), RPC (B. H. Sigelman, L. A. Barroso, M. Burrows, P. Stephenson, M. Plakal, D. Beaver, S. Jaspan, and C. Shanbhag (2010); S. Luo, H. Xu, C. Lu, K. Ye, G. Xu, L. Zhang, Y. Ding, J. He, and C. Xu (2021); Y. Gan, Y. Zhang, K. Hu, D. Cheng, Y. He, M. Pancholi, and C. Delimitrou (2019); 21)) and flow control (e.g., key-value store (Nishtala et al., 2013), web (98; 99)).
It is not obvious how to identify causal pairs. At any given time, many concurrent requests may be in flight. Connection persistence and stream multiplexing is standardized and widely deployed in the HTTP protocols (48; 49; 47). Two consecutive requests arriving at a vantage point from the same connection do not necessarily form a causal pair.
Our second key idea is to leverage the time gaps between packet arrivals to identify causal pairs. Once a client has sent a few requests, subsequent requests are typically blocked by dependencies or flow control. Hence, causally triggered requests arrive after a time gap longer than the inter-arrival times of the packets just prior. We call this the prominent packet gap assumption. By choosing an appropriate threshold for packet inter-arrival times, a vantage point can classify any packet arriving with a time gap exceeding this threshold as a causally-triggered request. The time interval between two consecutive causally-triggered requests provides one estimate of the response latency.
But what should the time threshold be set to? How can we make this setting robust across deployments (e.g., wide-area vs. data center), network load, transport dynamics (e.g., TCP flavors), and applications (e.g.. web vs. key-value store)?
Our third key idea is the following: While a single prominent packet gap is subject to (unknown) network, application, and server conditions, observing packet inter-arrivals over a period of time provides a more robust picture. In this paper, we leverage the probability distribution of inter-arrival times to identify prominent packet gaps by designing a lightweight construction to measure a coarse histogram of all packet inter-arrival times within a connection. We also devise a procedure to estimate the average response latency over the (configurable) time period when the distribution was measured.
We call our algorithm, a synthesis of the three ideas above, Pirate. We show how to use measurements from Pirate in a feedback control loop to adapt a layer-4 load balancer. The load balancer may use direct server return (Patel et al., 2013) to avoid processing the responses. Since Pirate works under routing asymmetry, it is possible to make latency-aware decisions by unilaterally changing the load balancer, leaving clients, servers, application software, and the network unmodified.
Results. In §4, we evaluate Pirate under a realistic web workload derived from the Alexa top 100 web sites, with a web server whose CPU availability varies according to a real CPU utilization trace (Verma et al., 2015). Across all monitored responses, Pirate achieves a median relative error of 0.63% relative to the response latency measured at the client application, and gets to within $\pm$15% error 90% of the time. Our results also show that techniques for measuring transport-layer RTT do not faithfully measure application-layer response latency, even when they can leverage bidirectional traffic visibility. We integrated Pirate into Katran (86), an open-source layer-4 load balancer based on Maglev hashing (Eisenbud et al., 2016). Latency-aware Katran cuts the 99th percentile latency in our experimental setup by 37% on average across loads, relative to vanilla Katran.
We outline caveats when interpreting our results in §5. A shorter version of this paper appeared previously at a workshop venue (citation elided for anonymity). In comparison, this paper incorporates new techniques to make estimation more robust, significantly expands evaluations of accuracy and overhead under realistic settings, and designs a feedback controller that operates on real-time measurements. This paper does not raise any ethical issues.
2. Motivation and Background
Monitoring latency is a fundamental requirement for the design, maintenance, and optimization of interactive Internet services. This paper studies the problem of measuring response latency, which we define as the time between when an application client sends out the request and when it receives the last byte of the corresponding response at the application layer. The client may be a user device contacting a web server or an RPC client in one tier of a distributed multi-tiered application.
2.1. Our Goals
We seek a passive measurement approach to continuously measure response latency at a vantage point that can lie outside the client and the server, along the path from the client to the server. To make it more broadly applicable, we also aim to achieve the following specific goals.
G1. Handle missing or obscured transport headers. Encryption of application-layer payloads is ubiquitous on the web and in compute clusters (1; 19; 81; 70). Modern network-layer security goes one level further, obscuring transport-layer headers. Deployments of network-layer encrypted tunnels (e.g., WireGuard (126), IPSec (51)) have been trending upward (116). Even when transport is unencrypted, overlay networks and IP fragmentation can cause transport headers to be missing from packets. We desire a method that does not require visibility into the transport-layer headers.
G2. Work under routing asymmetry. Nodes conducting passive measurement may be at locations that do not have access to traffic in both directions between the client and the server. The prevalence of routing asymmetry is well established in the wide-area Internet (Paxson, 1997) and also in data centers (Zhang et al., 2018). Further, many network deployments use asymmetric routing by design. For example, layer-4 load balancers employ direct server return (Patel et al., 2013; Eisenbud et al., 2016), bypassing the load balancer for response packets. BGP policies on the Internet (e.g., stub networks) may assign different ingress and egress border routers to a given connection (Katz-Bassett et al., 2010; Uhlig and Bonaventure, 2004).
G3. Generalize across transport dynamics. The design and deployment of new transport algorithms for congestion control and loss recovery is an active and rapidly evolving area (Ferreira et al., 2024). We seek techniques that avoid relying on specific transport-protocol behaviors, and sidestep the need for new transport-protocols extensions.
G4. Efficient enough to run in software middleboxes. Measurement and monitoring devices are frequently deployed as virtual network functions managed through a software-defined measurement infrastructure, e.g., (Palo Alto (2025); 117). We aim for techniques that impose low compute and memory overheads in software deployments, e.g., middleboxes running high-speed packet processing.
2.2. Prior Work and Its Applicability
In the academic literature, the problem most closely related to passively estimating response latency is the passive measurement of round-trip times (RTTs) at the transport layer.
RTT estimation approaches. The basic idea of passive RTT estimation is to relate data packets to their corresponding transport-layer acknowledgments (ACKs) and measure the time difference between the pair. Many prior works on RTT estimation determine this relationship by associating the sequence numbers or timestamps in data packets with the corresponding information echoed in ACK packets (Shawn Ostermann, ; Ghasemi et al., 2017; Sengupta et al., 2022; Ding and Rabinovich, 2015; Veal et al., 2005; Jaiswal et al., 2004). Implementing this basic approach requires visibility into transport headers in the clear (G1) and seeing packets in both directions (G2).
We are aware of only a small number of techniques that apply when the passive observer sees packet traffic in only one direction. An early approach by Jiang and Dovrolis (Jiang and Dovrolis, 2002) estimates RTT as the time interval between the SYN and the ACK packet of the 3-way handshake, and also between small packet bursts during the early rounds of slow start. These techniques are customized to the specific dynamics of the TCP protocol (G3) and do not measure connections beyond the initial phase. Another set of prior works estimate RTTs by computing frequency spectra from the time series of packet arrivals (Carra et al., 2010; Veal et al., 2005). Such computations require significant memory to hold packet timings (G4), and can be complex to tune (Ding and Rabinovich, 2015). Most recently, researchers have proposed the spin bit as a protocol extension (G3) to measure RTTs (De Vaere et al., 2018) passively. It has been incorporated as an optional mechanism in QUIC (115), but its deployment may be limited (Kunze et al., 2023).
Response latency differs from RTT. A high RTT implies a high response latency, but a low RTT does not necessarily mean a low response latency. In principle, the response latency can differ significantly from the transport-layer RTT. The server can send a transport-layer ACK to the client before sending any response data. Further, responses, typically larger than requests, may span multiple packets, widening the gap between the transport-layer ACK and the last byte of the response. When the client and server use a protocol with pipelined requests (e.g., HTTP/1.1), multiple application requests may be transmitted in a single packet from the client to the server. The transport-layer ACK only corresponds to the first response.
Figure 1. Transport round-trip times (RTTs) do not match response latency, with clients running (a) TCP HTTP/1.1 and (b) UDP QUIC. See discussion in §$2.$
Experimentally, we find that RTT and response latency differ even when there is only one small application-layer request in flight. We ran a web benchmark where a TCP client maintains a single outstanding request at a time with the server. Our full experimental setup is described in §4.1. On one client thread, Figure 1 (a) shows the time series of response latencies (marked req-to-res) against three techniques to measure transport RTT: (i) reading statistics from the kernel network stack through tcp-probe (113); (ii) estimating RTT from a packet trace (tcptrace (Shawn Ostermann, )); and (iii) syn-ack (Jiang and Dovrolis, 2002). Not only is the response latency different from the RTT, but it is also uncorrelated with it.
We ran a similar experiment to compare the RTT of a QUIC connection measured using the spin protocol extension (De Vaere et al., 2018) with the application-level response latency, measured with quic library lsquic (71) on a real web page. Figure 1 (b) shows that the two values diverge, especially when there are multiple RTT samples corresponding to a single response latency. The error bars show the minimum and maximum RTT reported through spin.
Other related work. Given the emerging prevalence of encrypted transports, some recent work aims to passively infer quality of experience metrics specifically for web applications (Huet et al., 2020; Wehner et al., 2021). These metrics compose the response latencies of multiple objects, e.g., to estimate overall web experience (87), through machine learning and prediction. In contrast, we are concerned with continuously measuring the response latencies of individual objects throughout the lifetime of a connection. There is also significant prior work on passively measuring the latency of specific network segments, e.g.(Kompella et al., 2009; Lee et al., 2010; Ghabashneh et al., 2022; Canel et al., 2024) and on algorithms to estimate latencies over end-to-end network paths using previously measured delays, e.g.(Ng and Zhang, 2002; Madhyastha et al., 2006). Plenty of active approaches exist to measure end-to-end network latency characteristics, e.g.(Sommers et al., 2007; Yu et al., 2015; Sundaresan et al., 2017). This paper seeks continuous passive measurement techniques for end-to-end response latency.
In summary, we believe that passive continuous measurement of response latencies under modern transport protocols and practical deployment constraints (§2.1) remains unsolved.
3. Pirate
This section introduces an algorithm to measure response latency passively and continuously for interactive applications (§2.1). We assume that the vantage point of measurement lies along the path from the client to the server.
3.1. Causal Pairs
Instead of measuring the time interval between a request and its response, our first key idea is to measure a proxy time interval—the time between a request and the packet transmitted by the client upon receiving the corresponding response. We call the latter packet a causally-triggered request, and the pair of packets a causal pair.
Causally-triggered requests exist due to several reasons: (1) Cross-request dependencies: Many interactive applications issue subsequent requests only when responses to previous requests have been received and processed by the client application. For example, web and RPC clients issue follow-up requests to fetch objects based on previous responses. (2) Flow control: Clients frequently cap the number of application or transport data that are outstanding at the server. Web browsers, memcached clients, and RPC client libraries are known to subject clients to such flow control (98; 99; 92). (3) Acknowledgments. A response may trigger an acknowledgment at the transport or application layer.
If a client application maintains exactly one request in flight, issuing the next request only after the response to the first arrives, then the two requests form a causal pair, and measuring the time interval between the two requests provides an estimate of the response latency. Causal pairs are a generalization of syn-ack estimation (Jiang and Dovrolis, 2002).
There are two caveats to using causal pairs to estimate response latency. First, a client application may take additional time to process a response before generating a follow-up request, due to considerations such as parsing or delays caused by process or thread scheduling. We call this the client think time. In experiments on lightly-loaded machines, we have measured client think times ranging from a few microseconds to a few tens of microseconds. Second, a response latency measurement is available only for one request per causal pair, not for all requests transmitted between the two requests of a causal pair on the same connection.
A key challenge arises when we go to settings where multiple requests or packets are concurrently in flight. Two consecutive requests observed at a vantage point from the same connection need not form a causal pair. Accurate knowledge of the client’s ongoing window size (at the transport or application layer) may help identify packets that are not causally related to each other, e.g., if they belong to the same window. However, inferring the window size typically itself requires assumptions on the dynamics of the transport protocol in question (Jaiswal et al., 2004; Ghasemi et al., 2017). This violates goal G32.1).
3.2. Prominent Packet Gap Assumption
Our second key idea is to leverage the timings of packet arrivals to identify causal pairs when a client has many concurrent requests in flight. The reasons that produce causally-triggered requests (cross-request dependencies, flow control, acknowledgments) result in clients transmitting a burst of data in one shot, then pausing request transmission until a response arrives. The sender side of transport often transmits in bursts to leverage hardware batching optimizations (e.g., TSO (112)), while receivers typically optimize I/O and CPU by batching ACK processing (112; 95). The prevalence of burst-then-pause behavior is well documented: TCP senders that use window-based transmission send bursts of packets, termed flowlets (Sinha et al., 2004; Vanini et al., 2017), separated by a pause, termed the flowlet gap.
We exploit the observation that, in many practical scenarios, the pause is noticeably longer and more prominent than the packet inter-arrival times in the burst, since the former is subject to network and server processing delays, while the latter is determined solely by the client’s ability to transmit requests. Hence, the first packet to arrive after the pause must be a causally-triggered request. The time interval between consecutive causally-triggered requests provides an estimate of the response latency.
The assumption of a prominent packet gap does not hold universally. Packets arriving at the vantage point may be uniformly paced either by the sender’s transport, e.g.(Cardwell et al., 2016; Goyal et al., 2022) or a bottleneck link or policer (Flach et al., 2016; Kanuparthy and Dovrolis, 2011). We conjecture that latency-sensitive applications transmitting small amounts of bursty traffic may not be widely subject to pacing, capacity bottlenecks, or policing.
Given a fixed time threshold $\delta$, packets that arrive with an inter-arrival time gap of more than $\delta$ are considered to be a part of different bursts. We estimate the response latency by the time between the first packets of successive bursts. This algorithm is shown in Algorithm 1 in Appendix A.
How should we choose the time threshold $\delta$? The threshold substantially impacts the accuracy of our estimate. Choosing a $\delta$ that is too large will miss legitimate causal pairs, only identifying long idle periods in the connection. Choosing a $\delta$ that is too small can make the estimation vulnerable to noisy gaps between packets transmitted within the “burst,” for example due to scheduling or processing delays at the client’s application or transport layer.
Fundamentally, the duration of the pause depends on several factors, such as the timing of writes from the client application into the transport layer, the client transport’s scheduling of packet transmission (i.e. flow and congestion control), scheduling at the network stack’s traffic control layer or the NIC (Saeed et al., 2017), cross traffic competing with the connection before packets arrive at the vantage point, the network round-trip time, and server processing delays. The combination of these factors makes it unlikely that a fixed threshold $\delta$ can work across different scenarios, or across time, even for a single connection in a specific network. Approaches that select time thresholds for flowlet-based network load balancing (Sinha et al., 2004; Kandula et al., 2007; Vanini et al., 2017; Alizadeh et al., 2014) do so either by using the differences in expected path latencies or by using an an empirically tuned value that achieves a desired balance among paths. These approaches are unsuitable for a passive observer attempting to measure latency.
3.3. Choosing a Packet Gap Threshold
Our third key idea is that observing the timing of packet arrivals over time can provide meaningful clues to a good value for the pause-time threshold $\delta$3.2). This observation is inspired by prior work that motivates a longitudinal view of noisy data for robust measurement and feedback control (Geng et al., 2018; Kelly et al., 2008; Wischik and McKeown, 2005; Raina et al., 2005). Instead of producing an estimate instantaneously, we can produce an average over a configurable time epoch.
Specifically, given an empirical distribution of inter-packet time gaps (IPGs) observed over the lifetime of a connection, the distribution’s significant modes carry information about specific events occurring on that connection. In Figure 2, we show the empirical distribution of IPGs in a simple experimental scenario where a TCP client has roughly constant response latency across all requests but experiences occasional packet losses and idle periods (when no data is transmitted or received). The IPG distribution includes a batch of packets within a burst (the smallest-valued modes); IPGs across bursts of causally dependent packets; loss timeouts (typically set larger than response latencies); and idle periods (the largest modes).
Figure 2. Modes in the empirical distribution of inter-packet gaps (IPGs). Modes carry useful information about phenomena of interest occurring over the measurement epoch.
Through knowledge of standard parameters used for retransmission timeouts in transport stacks (publicly known for several standard TCP and QUIC configurations (94; 20; V. Vasudevan, A. Phanishayee, H. Shah, E. Krevat, D. G. Andersen, G. R. Ganger, G. A. Gibson, and B. Mueller (2009))), it is possible to eliminate the modes corresponding to retransmission timeouts and idle periods, leaving behind the de-noised distribution of IPGs, which only includes gaps between request packets that are either (i) back-to-back transmissions without causal dependencies, or (ii) causally-triggered requests sent after a prominent packet gap.
[…]
Our use of IPG distributions differs from prior techniques that probe the network using packet trains and use IPG distributions to estimate bottleneck link bandwidth, e.g.(Dovrolis et al., 2001; Prasad et al., 2003). The average response latency is the sum of multiple packet gaps; identifying which packet gaps to combine and in what proportion is the core challenge of latency measurement, a challenge prior work does not address.
[…]
3.4. Maintaining Efficient Histograms
To maintain an IPG distribution with memory efficiency, we design an algorithm that maintains a small number of buckets by dynamically varying the resolution of each bucket according to the IPGs observed on that connection. Dynamic bucket resolution has the advantages that (1) the histograms of different connections may freely span different ranges of IPGs; (2) we can avoid the overhead of bucket counters for IPG values that do not occur; and instead, (3) focus the available memory on IPG values that are actually observed.
Our algorithm works as follows. The histogram, $M$, includes a (configurable) maximum number of buckets (modes) $N$. Initially, all buckets are uninitialized. Each bucket (after initialization) specifies the minimum, maximum, count, and sum of all the IPGs observed within that bucket. For each observed IPG $g$, the algorithm either (i) counts $g$ into an existing bucket whose (min, max) includes $g$, or (ii) if $g$ is “close enough” to an existing bucket, extends the bucket or merges two buckets to produce a new, larger bucket that now counts $g$, or (iii) puts $g$ into its own new bucket if there is an available uninitialized bucket, or (iv) discards $g$. As few as $N=10$ buckets (maximum number of modes) per connection prove to be sufficient to capture all IPG distributions in our experiments. At the end of each epoch, the proportional mode sum (§2) is computed over the average mode values of $M$ to emit a response latency averaged over the epoch. Our complete algorithm is shown in Algorithm 2 in Appendix B.
Similar to ours, there exist storage-efficient algorithms in the streaming setting, that use a dynamic bucket size to maintain modes efficiently, e.g.(Mitzenmacher et al., 2012). We leave the adaptation of such algorithms to our scenario and a quantitative comparison to future work, since our experiments (§4) show that our current algorithm is practical.
3.5. A Latency-Aware DSR Layer-4 Load Balancer
To showcase the utility of real-time response latency measurement under asymmetric routing, we show the design of a latency-optimizing layer-4 load balancer. Layer-4 load balancers implement direct server return (DSR). This mechanism allows backend servers to send responses directly to the client, bypassing the load balancer (Patel et al., 2013) along the path from the server to the client. To our knowledge, this work is the first to enable the availability of real-time latency signals at DSR load balancers unilaterally without modifying the client, server, application, or the network. We do not claim any novelty in our algorithm design; there is a significant body of algorithmic work on performance-aware request load balancing, e.g.(L. Suresh, M. Canini, S. Schmid, and A. Feldmann (2015); B. Wydrowski, R. Kleinberg, S. M. Rumble, and A. Archer (2024); 46; T. Barbette, C. Tang, H. Yao, D. Kostić, G. Q. M. Jr., P. Papadimitratos, and M. Chiesa (2020); M. Kogias, G. Prekas, A. Ghosn, J. Fietz, and E. Bugnion (2019); M. Mitzenmacher (2001); R. Gandhi and S. Narayana (2025); 46). Below, we describe one way to design a feedback controller based on latency measurements.
Our design augments a Maglev-hashing-based load balancer (Eisenbud et al., 2016). We assume that the load balancer provides mechanisms to assign weights to distribute the load across servers. We use three key ideas to assign the weights. First, we take away weights from servers with latencies above the high watermark, and place them on servers with latencies below the low watermark. Second, we limit adding weight to servers with insufficient recency (“freshness”) in their latency measurements. Third, we regress to the mean by slowly equalizing weights across servers when latencies have not changed recently. Appendix C presents a complete discussion of these ideas and our design.
4. Evaluation
In this section, we empirically evaluate Pirate under the experimental settings described in §4.1. We ask the following questions:
4.2) How accurate is Pirate in measuring application-level response latencies under realistic applications and settings? How does Pirate compare to transport RTT estimators?
4.3) What overheads does Pirate impose on a software middlebox?
4.4) Can measurements from Pirate enable real-time robust feedback control?
4.5) How well does each core idea from Pirate work in realistic settings?
4.6) Is Pirate robust to network and load variability?
4.1. Experimental Setup
Implementation of Pirate. Pirate runs as a kernel bypass program developed in Linux eBPF, which attaches to the express data path (XDP (Høiland-Jørgensen et al., 2018)) hook, which resides in the network device driver in the kernel. We use a standalone XDP forwarder that implements the Pirate algorithm for the accuracy experiments. For our evaluation of the feedback controller, we instrumented the Katran layer-4 load balancer (Shirokov, 2018), developed and open sourced by Meta. Pirate runs as a program that calls Katran using a BPF tail call. Katran implements direct server return (Patel et al., 2013). The measurement component of Pirate comprises approximately 500 lines of C code.
Web benchmarking framework. We use the WebPolygraph suite (123) as our HTTP/1.1 client and server for benchmarking. WebPolygraph allows detailed configuration of workload characteristics such as (1) the number of benchmarking threads, offered load (requests per second) initiated from each benchmarking client thread, and the number of concurrent requests on persistent pipelined HTTP connections over TCP; (2) the types of objects (e.g., jpeg, etc.) and their prevalence (e.g., 20%) among the requested objects; (3) properties of the dependency tree of objects requested starting from the root URL requested by the client (e.g., an object of type $T$ includes $N_{a}$ objects of type $a$, $N_{b}$ objects of type $b$, and so on, where the values of the $N_{(.)}$ can be drawn from a custom probability distribution; and (4) the probability distributions of the sizes of each object type. The WebPolygraph server is single-threaded. With our client workload (see below) and server machine, each client-to-server connection saturates a single CPU core at a load of 4K requests/second.
Client workload. We populate web page parameters (types, sizes, object dependencies) into WebPolygraph by creating a dataset of web objects from the home pages of a random sample of 20 unique web sites from the Alexa top-100. Each HTTP request issued by the client fetches one of 4 kinds of objects (HTML, CSS, JS, image), with the response size drawn from the empirical size distribution of the corresponding object type in our dataset (CSS: median 11.8 KByte, max 1.2 MByte; JS: median 19.6 KByte, max 4.05 MByte; HTML: median 39.5 KByte, max 2.65 MByte; image: median 20.8 KByte, max 8.86 MByte). Unless specified otherwise, our client runs on single machine with 20 threads and 200 total connections, driving an aggregate workload of 2K requests/sec. To simulate dependent HTTP requests generated by parsing HTTP responses (e.g., images embedded in HTML), each response returned by the server includes references to different kinds of embedded objects drawn from the corresponding empirical probability distribution in our dataset. Connections are persistent and pipelined. Unless stated otherwise, each WebPolygraph client connection maintains at most 4 outstanding requests, comparable to default limits in the HTTP standard and in popular browsers (47; 18; 83; 97; 32). Each active connection may issue 1–7000 requests, with a median of 2094 HTTP objects per connection.
Server performance variability. We derive a time series of CPU allocations for our benchmark web server from a real CPU utilization trace of Google’s Borg cluster jobs (A. Verma, L. Pedrosa, M. R. Korupolu, D. Oppenheimer, E. Tune, and J. Wilkes (2015); 41) (2019 trace). The trace provides histograms of CPU usage over a 5-minute period, where each sample collected in the original data is over one second. We select the time series of the job with the largest average CPU utilization in the trace, sample from the histograms, and enforce the resulting CPU allocation on the WebPolygraph server using Linux cgroups.
[…]
Figure 3. The experimental topologies we use on CloudLab (Duplyakin et al., 2019). Most experiments use the single cluster “triangle” topology (a). We also run experiments on a wide-area topology (b) and a star topology (c). Traffic flows from the clients (left) to servers (right) via the measurement vantage point (magnifying glass). Server to client traffic bypasses the vantage point. Link bandwidths (Mbit/s) are shown.
Topology. We set up our implementation of Pirate and the client/server workloads on CloudLab (Duplyakin et al., 2019).
[…]
We also evaluate Pirate in wide-area settings with the topology shown in Figure 3 (b). A client machine and a contending traffic source reside in the CloudLab Utah site and send traffic over a bottlenecked wide-area link to the Clemson site, which runs the measurement vantage point. The corresponding servers reside at the Wisconsin site. Servers return traffic directly to the client, bypassing the measurement vantage point. This network exhibits a propagation round-trip time of 70 milliseconds. Link bandwidths (Mbit/s) are labeled in the figure. The bandwidths along the forward and reverse paths are asymmetric.
To test load balancing, we also employ a star topology (Figure 3 (c)), with two clients connecting to four servers via a measurement vantage point. The seven server machines include an Intel E5-2640v4 10-core CPU, 64GB memory, and dual-port 25Gb NIC. A Dell S4048 switch interconnects the machines. We configured the switch and all the servers to forward clients-to-servers packets via the vantage point, and server-to-client packets directly via the switch, mimicking a direct server return configuration (57). The Pirate algorithm runs on the switch-facing ingress interface of the vantage point.
Techniques compared.
(1) Ground truth response latency (req-to-res): We instrumented the WebPolygraph client to compute the time delay between the transmission of each request and the reception of the complete response object for that request at the application layer. This is our ground truth for response latency, labeled req-to-res on the graphs.
(2) Causal pair delay (req-to-req): Even when the reception of a complete response could trigger another request (i.e., a causally-triggered request), a client application may incur additional time to transmit the latter, due to processing (e.g., parsing) and scheduling delays. Since Pirate approximates the time delay between the causal pair, for reference, we also show the time between the request and the triggered request at the client, labeled req-to-req on the graphs.
(3) Transport RTT measurement: We compare Pirate against transport-level RTT estimators, specifically (1) tcptrace (Shawn Ostermann, ), an open-source tool that analyzes pcap traces, (2) tcp_probe (113), which emits the sample RTTs maintained by the Linux network stack; and (3) syn-ack delay (Jiang and Dovrolis, 2002; Szymaniak et al., 2008), the time between the SYN and the ACK packets in the TCP 3-way handshake.
Syn-ack only produces one measurement per connection. We ensure that the values reported by the aforementioned techniques are directly comparable by aligning the transport-layer sequence numbers represented in the measurements. All techniques except Pirate and syn-ack can see traffic in both directions.
Accuracy metrics. When we compare a measurement technique $T$ against a baseline technique $B$, we report the absolute error $latency_{B}-latency_{T}$ and the relative error $(latency_{B}-latency_{T})/latency_{B}$ as a percentage.
4.2. Accuracy of Latency Estimation
[…]
Figure 4. Accuracy in estimating the response latency (req-to-res) under realistic settings. Subfigures (a)–(c) show distributions within a single cluster; (d) shows Pirate’s relative error in a wide-area network with different degrees of cross traffic.
How close does Pirate get to the ground truth response latency observed at the client application (req-to-res)?
Single-cluster setting. In a triangle network topology (Figure 3 (a)), our client sends a total of 2K requests/second to the server across all threads and connections.
Figure 4 (a) shows the CDF of the response latencies at the client (req-to-res) and the vantage point (Pirate) across all objects over all connections. We also show the time delay between the causal pair (req-to-req), which Pirate approximates. The three distributions are closely aligned, providing confidence that (i) the time delay between causal pairs is a good estimator of response latency, and (ii) that Pirate can estimate the latter with high accuracy, in realistic settings where the client maintains persistent and pipelined connections with multiple concurrent application-layer requests over each connection (§4.1).
Figure 4 (b) shows the CDF of the absolute error of Pirate and req-to-req against req-to-res, while Figure 4 (c) shows the CDF of the relative errors. The causal pair delay is always an overestimate of the response latency, showing a relative error between $10\%$ to $0\%$. Pirate’s relative error is less than 1% at the median, but stretches beyond $\pm$10% outside $10^{th}-90^{th}$ pc, and beyond $\pm$15% at the tails beyond $5^{th}-95^{th}$ pc. In a state of the art where response latency measurement is nonexistent under asymmetric routing, an estimator that gets within a 15% error 90% of the time is useful (§3.5; §4.4).
[…]
Figure 5. Classifying Pirate’s errors in estimating response latency, showing prevalence (a) and magnitude (b) across connections. Discussion in §4.2 (why does Pirate make errors?)
Why does Pirate make errors? An investigation of when and how significantly Pirate miscomputes response latencies in the experiment above reveals two kinds of errors. Figure 5 measures their prevalence and magnitude.
The first kind of error, labeled parsing-scheduling, arises when a client has a request to transmit, and also has a chance to transmit (i.e., not subject to flow control or pipeline concurrency limit), and yet takes additional time to transmit. This is due to response processing (i.e., parsing) and delays in scheduling the client process. Pirate correctly identifies the causal pairs but the latter delay is not representative of response latency.
The second kind of error, labeled staggered-response, occurs when a series of large response objects (each spanning multiple network RTTs) causes successive requests to be transmitted well apart in time. Pirate mistakes the pauses between the causally-unrelated successive requests to be pauses between causally-related ones. Here, Pirate mis-identifies the causal pairs in the first place.
Figure 5 (a) shows the CDF of the fraction of erroneous samples per connection. The parsing-scheduling error is more prevalent. Roughly 70% of the connections have fewer than 5% erroneous samples of either kind. Figure 5 (b) shows the CDF of the average relative error of the erroneous samples per connection. The magnitude of these errors is consistent with the overall distribution (Figure 4 (c)). Further, the magnitudes of the two kinds of errors are comparable.
Wide-area setting. We run the client workload from §4.1 over a wide-area network spanning three sites (Figure 3 (b)). The network includes cross traffic contending with the measured client-server connection. The cross traffic consists of a mix of short- and long-lived TCP flows whose sizes are sampled from a distribution measured in a production network at Microsoft (Alizadeh et al., 2010). Flows arrive and depart throughout the lifetime of the experiment. The arrival times of new flows are sampled from a Poisson distribution parameterized by the average aggregate rate of cross traffic, a value we configure as a fraction of the bottleneck link capacity of the client-to-server path. Figure 4 (d) shows the relative error of Pirate across different levels of cross traffic, expressed as a % of bottleneck capacity. Pirate’s error is close to 0% for more than 90% of the measurements across all cross traffic configurations. Comparing Figure 4 (c) and Figure 4 (d), Pirate exhibits a higher accuracy in larger RTT, wide-area settings, than in smaller RTT, single-cluster settings.
[…]
Figure 6. (a) CDF of all measurements throughout the lifetime of all connections. (b) and (c): Time evolution of the latency of a single (randomly chosen) connection during a small window of data transfer.
Comparison Against RTT estimators. To perform a faithful comparison of Pirate against transport-layer RTT estimators (§2.2), we run a simplified web workload in which clients from §4.1 hold only a single request in flight per connection. A majority of response objects in the workload are small enough to fit into a single packet, which increases the likelihood that transport RTTs can faithfully model response latency. We use the single cluster setup (Figure 3 (a)). All compared techniques except syn-ack and Pirate (tcp-probe, tcptrace, req-to-req, and req-to-res) have visibility into both directions of network traffic.
Figure 6 (a) compares the CDF of the latencies measured by each technique over all objects across all connections. Even with a simplified version of a realistic workload, RTT measures consistently underestimate response latency, since they are more closely aligned with the time to the first byte of the response, as opposed to the last byte. Figure 6 (b) and (c) show measurements reported over a small interval of data transfer. Syn-ack estimation produces only one measurement per connection, which does not capture latency variability over the lifetime of the connection.
[…]
Figure 7. Latencies from memtier (ground truth) and Pirate.
Measuring API traffic from memcached. To evaluate the efficacy of Pirate for API traffic, we run a memcached benchmarking client, memtier (76) in the single-cluster topology (Figure 3 (a)). The client runs with 10 threads, each with one client connection. The proportion of get and set requests is 1:1. The response data size is configured to be in the range 40–10000 bytes. The benchmark was run for 2 minutes. Figure 7 shows the CDF of the response latency measured at the application layer by memtier (ground truth) against estimates from Pirate. Moving beyond web workloads, Pirate may be useful to estimate response latencies in delay-sensitive applications more generally, when such applications exhibit flow control or request-level dependencies.
4.3. Software Overheads
[…]
Figure 8. The CPU and latency overheads of Pirate in the context of a layer-4 load balancer. See discussion in §4.3.
We evaluate the CPU overheads of the Pirate algorithm and its XDP implementation in the context of another XDP packet-processing application, Katran (86), a layer-4 load balancer used in production at Meta. In our single-cluster setup (Figure 3 (a)), the vantage point serves as the device under test, running Pirate and Katran chained via an eBPF tail call (28). The client runs wrk2, a web benchmarking client, with 4 threads and a configurable offered load (requests/sec) and connection count. For simplicity of interpretation, we direct all requests to a single fixed CPU core using Linux Receive Side Scaling (RSS (101)).
In Figure 8 (a) and (b), we fix the number of connections to 64 and vary the offered load (Krequests/sec). Subfigure (a) reports the CPU utilization measured using perf events (38), in comparison to Katran and the CPU idle thread. Pirate’s CPU usage is comparable to Katran’s but slightly higher. Figure 8 (b) shows the time per packet across loads for Pirate. To put this in context, on our experimental system, Katran incurs 1115 ns per packet, while a simple forwarding/redirect program incurs 870 ns.
In Figure 8 (c) and (d), we fix the total offered load at 8000 requests/sec and vary the number of active client connections. Subfigure (c) shows that Pirate offers stable packet-processing times and throughput upto 128 connections, but degrades beyond this point. An investigation using system performance counters (38), reported in subfigure (d), shows that the branch miss and data cache miss rates rise sharply beyond 128 additional connections.
The results above have been measured with all packet processing executed on a single CPU core. Practical systems scale single-core throughput and connections nearly linearly with additional CPU cores (101; T. Barbette, G. P. Katsikas, G. Q. Maguire, and D. Kostić (2019); F. Pereira, F. M.V. Ramos, and L. Pedrosa (2024); Q. Xu, S. Miano, X. Gao, T. Wang, A. Murugadass, S. Zhang, A. Sivaraman, G. Antichi, and S. Narayana (2025)). Hence, the total throughput and the number of active connections supported in a real deployment would depend on the available CPU parallelism.
XDP requires provisioning memory in advance for lookup data structures. By default, Katran provisions a hash table for 64K connections, incurring 458 MBytes. In comparison, supporting the same number of connections in Pirate requires 12 MBytes (154 bytes/connection).
4.4. Feedback Control
Figure 9. Comparison of 99th percentile tail latency of an unmodified Katran load balancer and our latency-aware one.
Can measurements from Pirate enable more robust real-time feedback control? To answer this question, we evaluate the use of measurements from Pirate within a layer-4 load balancer, Katran (86), implementing direct server return (Patel et al., 2013), using the approach discussed in §3.5. We use the star topology in Figure 3 (c), where two clients connect to four servers via a vantage point that uses the measurements to balance server load. We simplified the web workload in §4.1 so that each client connection maintain only one request in flight. Figure 9 compares the 99th percentile tail response latency of the latency-aware load balancer with that of the vanilla version at varying offered loads (Krequests/sec). Latency awareness produces, on average, a 37% reduction in tail latency across loads. Moreover, latency awareness also results in more predictable tail latencies (smaller error bars).
4.5. Benefits of Key Ideas
[…]
Figure 10. Viability of the approximations in Pirate. See discussion in §4.5. The graphs show the relative errors of (a) req-to-req over req-to-res; (ii) prominent-gap over req-to-req; and (iii) Pirate over prominent-gap, at three pipeline depths.
We evaluate the viability of the core ideas in Pirate through microbenchmarks. Here are the key questions we are interested in: (i) Does the time gap between causally-related packets reflect the application-level response latency (§3.1)? (ii) How likely is a packet following a prominent time gap (i.e., a time gap “significantly” larger than other inter-packet time gaps) to be a causally-triggered request (§3.2)? (iii) Are prominent time gaps captured accurately by maintaining distributions in Pirate3.3)?
We run our web benchmarking client and server (§4.1) within a single cluster (Figure 3 (a)). To simplify the experiment, we leave server performance unconstrained by interfering CPU allocations (§4.1), and collect measurements over one long-lived client-server connection. We measure: (1) the ground truth response latency (req-to-res), and also three successively coarse approximations of it: (2) causal pair delay (req-to-req, §3.1); (3) an approach that uses a threshold of 0.6 $\times$ the true response latency, to separate packets into batches of causally-related packets (§3.2), labeled prominent-gap; and (4) Pirate3.3). We align the transport-level sequence numbers reported across techniques in the measurements, so that estimates can be compared directly against each other. In our experiments, the number of concurrent requests in flight is a key factor that impacted measurement accuracy, hence we evaluate three pipeline depths (4, 8, 16).
Figure 10 shows the CDF of relative error of each approximate method above with respect to the method before it. Subfigure (a) shows that req-to-req is a good overapproximation of req-to-res. Subfigure (b) shows that prominent-gap is slightly erroneous but is still a good approximation of req-to-req. Subfigure (c) shows that Pirate is a similarly good approximation of prominent-gap. We also note the errors in the tail worsen as pipeline depth increases.
4.6. Robustness
[…]
Figure 11. (a) CDF of latencies for a loss rate of 1% (b) CDF of latencies for a packet reorder rate of 25%
Packet Loss and Reordering. We evaluate the impact of packet loss and reordering, both of which affect the packet inter-arrival times due to transport adaptation, on the Pirate’s estimation. In our single-cluster setup (Figure 3 (a)), we induce loss and reordering at varying rates over packets from the client towards the vantage point. (Loss on the server-to-client path triggers retransmission from the server to the client, which is not observed by the vantage point, hence we do not evaluate this.) Figure 11 (a) shows the CDF of latencies measured by Pirate at a loss rate of 1%, while Figure 11 (b) shows the CDF of latencies at a high packet reorder rate of 25%. Pirate produces robust estimates of response latency under both conditions. The estimation accuracies were similar across other loss and reorder rates we evaluated.
[…]
Figure 12. (a) Average response latency across all connections vs. offered load. (b) Relative error at 12K requests/sec.
Robustness to Offered Load. We evaluated the accuracy of latency estimation as the load offered to the server varies. In this experiment, we simplify the web workload from §4.1 to transmit only one request in flight, and vary the offered server load (requests/sec). Figure 12 (a) shows the average latency across all connections across load. At a load of 12K requests/second, Figure 12 (b) shows the CDF of relative error (averaged per connection) across the techniques. Pirate shows strong agreement with the response latency.
5. Discussion
This paper presented Pirate, an algorithm that passively and continuously measures response latencies under routing asymmetry, even when transport headers are missing due to encryption or fragmentation. Pirate leverages the idea of causal pairs, two requests the second of which is triggered by the response to the first. In realistic experiments, Pirate shows high accuracy and enables robust feedback control.
Pirate has two significant limitations, both stemming from the estimator’s assumptions. First, applications with significant client think times suffer from sizable overestimations of response latency (e.g. DASH video streams). Second, the estimator assumes that successive batches of packets are causally related. However, large response objects (§4.2), and arbitrary reordering in server responses (relative to requests) both violate this assumption. The latter challenges generalizing our results to HTTP/2 and QUIC.
Despite these limitations, Pirate is robust for HTTP/1.1-like web and fast-growing API traffic (65; 30), in the presence of multiple concurrent pipelined requests, server fluctuations, and network variability. These results may have broader implications for the design of reactive feedback control systems that rely on continuous passive measurement.
Appendix A An Estimator based on Prominent Packet Gaps
See Algorithm 1.
1:Fixed threshold on inter-packet gaps, $\delta$
2:Timestamp of the current packet’s arrival, nownow
3:The last time a new batch arrived for flow $f$, f.time_last_batchf.time\_last\_batch
4:The last time a packet arrived for flow $f$, f.time_last_pktf.time\_last\_pkt
5:An estimate of flow $f$’s reponse latency, $\hat{T}_{LB}$, if a new sample is produced, else undefundef
[…]
7:if $now-f.time\_last\_pkt>\delta$ then $\triangleright$ New batch: record response latency
8:  $\hat{T}_{LB}\leftarrow now-f.time\_last\_batch$
9:  $f.time\_last\_batch\leftarrow now$
10:end if
[…]
12:return $\hat{T}_{LB}$
Algorithm 1 Track causally-triggered requests using a fixed time threshold $\delta$ at the vantage point. The algorithm is executed upon receiving each packet of a flow $f$.
Appendix B Maintaining Efficient Histograms
See Algorithm 2.
1:New observation of an IPG $g$
2:A representation of the empirical probability distribution, $M$, with NN modes. For each $1\leq i\leq N$, the $i^{th}$ mode is a tuple $(min,max,count,sum)$, denoting the minimum, maximum, count, and sum of all the IPGs observed within the mode.
3:$M$ is updated with the additional IPG $g$
4:function UpdateModes(IPG $g$)
5:  for $m\in M.get\_modes():$ do
[…]
7:   left=m.get_min()left=m.get\_min()
8:   right=m.get_max()right=m.get\_max()
9:   if $left-\epsilon\leq g\leq right+\epsilon$ then
[…]
11:     AddGapToMode($g,m$)
12:     if $left-\epsilon\leq g\leq left$ then
[…]
14:      m.set_min(g)m.set\_min(g)
15:      ConsiderMergeModes(m.get_prev(),$m$.get\_prev(),m)
16:     else if $right\leq g\leq right+\epsilon$ then
[…]
18:      m.set_max(g)m.set\_max(g)
19:      ConsiderMergeModes(m,m.get_next()m,m.get\_next())
20:     end if
21:     return
22:   end if
23:  end for
24:  if $M$ has fewer than NN initialized modes then
[…]
26:   AddMode(M,g)(M,g)
27:  else
28:   discarded += 1
29:  end if
30:end function
Algorithm 2 Maintaining a small number of modes NN from the empirical probability distribution of IPGs.
Appendix C Detailed Design of the Latency-Aware Load Balancer
Below, we present the details of the design of our latency-aware layer-4 load balancer, which builds on a Maglev-like load balancer (Eisenbud et al., 2016) which permits the setting of weights to servers. The load balancer should measure response latency for a number of connections mapped to each server using the algorithms in §3, to produce a representative average latency for each server. We then use these average server latencies to adapt the weights assigned to the servers, using three key ideas.
First, we take away weights from servers which have average latencies larger than a high watermark, and place them on servers which have average latencies smaller than a low watermark. Our high (respectively, low) latency watermark is defined as $\alpha_{high}$ (respectively, $\alpha_{low}$) times the latency of the server with the smallest average latency. We use $\alpha_{high}=1.5$ and $\alpha_{low}=1.2$. Moreover, the weight shifted from a high-latency server is proportional to its latency.
Second, we restrict the low-latency servers eligible for placing additional weight by limiting ourselves to servers that have sufficient freshness in their latency measurement. As prior work has observed (Wydrowski et al., 2024), measured latencies are a good metric for the past performance of backend servers, but not their future. Instead, a different metric such as the number of requests in flight to a server is a leading indicator for the future performance of that server. Consequently, some prior works consider a combination of requests in flight and latency to balance load (Wydrowski et al., 2024; Suresh et al., 2015). A DSR load balancer cannot directly measure the number of requests in flight.
Instead, we measure the recency of the latency measurement of servers, by defining the freshness of a server latency measurement to be the ratio of the total number of requests received in the last measurement interval to the number of concurrent active connections at the end of that interval. This metric captures the intuition that a server on the verge of slowing down will have processed fewer requests per active connection. Conversely, a fast server must have processed more requests in the last measurement interval even if several of those connections have arrived and completed. We only consider a low-latency server eligible to take on additional weight if its freshness is at least as high as any high-latency server. Each such server receives an equal share of the total weight that is shifted away from the high-latency servers, subject to a cap on the per-server increment in weight, to avoid the thundering herd problem (Nishtala et al., 2013).
The third key idea is to regress to the mean: if latencies have not changed in the recent $k$ measurement intervals, the weights are slowly equalized across servers (we use $k=3$). Server and network performance can be highly variable. It is faster to improve performance from an operating point where no one server is assigned a disproportionately large weight.
Appendix D Heuristics in Pirate
To further improve the accuracy of Algorithm 2, we use two heuristics that eliminate IPG modes that we deem noisy. First, we coalesce IPGs corresponding to pure transport-layer acknowledgments into one IPG, by ignoring them from the stream of IPGs observed for a connection. The intuition is that pure ACKs do not represent the completion of an application-layer response, but rather signal partial completion. Mechanically, when transport headers are unencrypted, it is easy to identify pure ACK packets by inspecting the transport-layer headers, for example, the TCP ACK flag and packet size. When the transport layer is encrypted, this heuristic only applies if there is information that helps classify a packet as a pure ACK, e.g. payload sizes. Our second heuristic is to explicitly mark IPGs following a non-MTU packet as candidate modes to represent the IBG used in the proportional mode sum computation. The intuition is that clients typically send full MTU packets whenever packets can be transmitted independently and back to back.
[…]
Figure 13. (a) Pirate vs Pirate without both heuristics. (b) Pirate vs the request-to-triggered request delay.
Using our setup described in §4.1, we evaluate the benefits of the ACK-coalescing and MTU size heuristics (§3.4) in improving the accuracy of Pirate’s estimation. Figure 13 (a) compares the estimated response latencies when turning off both heuristics, against Pirate. In (b), the time series of request-to-triggered-request delays is shown against Pirate for reference. The heuristics provide a noticeable improvement to accuracy, especially in allowing Pirate to track fast fluctuations in latency.